OpenStudy (anonymous):

Prove that this sequence converges: s=1+1/2+1/3+...+1/n-ln(n)

OpenStudy (anonymous):

\[S=1+(1/2)+(1/3)+...+(1/n)-\ln(n)\]

OpenStudy (amistre64):

isnt that a series tho?

terenzreignz (terenzreignz):

It is :D But the series can't converge if the sequence doesn't converge, right? :D

OpenStudy (amistre64):

those are state secrets that must never be revealed .... or at least they are details that I never remember :)

OpenStudy (amistre64):

i know 1/n is divergent on it own; but that doesnt help

terenzreignz (terenzreignz):

@amistre64 that, and I don't know what "ln" is doing in this series, it seems out of place... is it though?

OpenStudy (amistre64):

it does seem rather "odd" lol

OpenStudy (anonymous):

Well, the serie should be \[(\sum_{k=1}^{n} 1/k) - \ln(n)\] doesn't?

terenzreignz (terenzreignz):

It will diverge... It's what you call the harmonic series, unless I'm missing something here What's n supposed to be? :)

terenzreignz (terenzreignz):

Wait a sec, this isn't an infinite series at all! Then doesn't that mean that for sure, the series will converge? :)

OpenStudy (anonymous):

That's the problem, it must converge, cause the limit of this sequence is Euler constant

OpenStudy (amistre64):

the limit as n to inf i believe

OpenStudy (amistre64):

can you define a rule for the partial sums? and then work that

terenzreignz (terenzreignz):

What is n, though? Is it a random number? Or perhaps, you let n approach infinity?

OpenStudy (amistre64):

s1 = 1 s2 = 3/2 - ln(2) s3 = 3/2 - ln(2) +1/3-ln(3) = 11/6 -ln(2/3) hmmm i think they forgot the limit on the front

OpenStudy (anonymous):

n takes the last value for n in the serie, as S=1+1/2+1/3-ln(3) when k goes from 1 to 3

OpenStudy (amistre64):

and -ln(n) = ln(1/n)

terenzreignz (terenzreignz):

Are you tasked to prove that it converges, or to prove that it converges to e?

OpenStudy (anonymous):

Prove that the sequence converges

OpenStudy (amistre64):

good luck, work beckons ....

OpenStudy (amistre64):

seq to me suggests that the seq of partial sums as n to inf

OpenStudy (anonymous):

Show that it's a cauchy sequence: \[S_{n-1}-S_n=ln(n)-ln(n-1)-\frac{1}{n}\rightarrow0\]

OpenStudy (anonymous):

Every real number cauchy sequence converges.

OpenStudy (blockcolder):

If I remember right, Sn is decreasing (or increasing?) and bounded, therefore, by Monotone Convergence Theorem, Sn converges.

OpenStudy (anonymous):

Yes, I was working on that, I prove it's increasing but what about the bounds?

terenzreignz (terenzreignz):

Doesn't the monotone convergence theorem only apply to the convergence of the sequence?

OpenStudy (blockcolder):

And the sequence in question is \(S_n=H_n-\ln(n)\), where Hn is the nth harmonic number. @knock I think a geometric point of view may help. Consider the function \(f(x)=\frac{1}{x}\).

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Why am I being ignored :p, I already proved it.

terenzreignz (terenzreignz):

@blockcolder I think only the (1/k) is in the sigma; ln (n) is subtracted from the sum, right?

OpenStudy (blockcolder):

Yeah, and...?

terenzreignz (terenzreignz):

and nothing I was just clarifying that ln(n) is not part of the sequence, only that it's subtracted from the series.

terenzreignz (terenzreignz):

The series does converge, for sure (since it's finite) When you say "its limit" what do you mean by it? Do you mean as n goes to infinity?

OpenStudy (anonymous):

The limit of this sequence as n goes to infinity is a number, called Euler constant and represented by leter gamma

OpenStudy (anonymous):

But I just want to prove that the sequence converges

OpenStudy (anonymous):

Ok here's a solution I found http://www.proofwiki.org/wiki/Euler-Maclaurin_Summation_Formula

terenzreignz (terenzreignz):

I think if all you want to do is prove convergence, then, the integral test should suffice?

OpenStudy (anonymous):

Yes but, as ln(n) it's not in the serie, can we use integral test?